# Transformation trees - documentation of multimodal image registration

### v2

@@ 2.2 Existing Approaches to Storing Image Transformations
Current methods for storing registration results vary depending on the system and standard used. However, none of them provide comprehensive documentation of all registration steps, as they typically focus only on storing the final transformed image or a single transformation matrix. This approach is found in tools such as MeVisLab [9], ITK [10], and Elastix [11], which allow image registration but often do not save the full history of transformations. These solutions enable quick access to a registered image but do not store details about the applied operations, making it impossible to reconstruct the registration process.
There are also several tools designed for registering triangular meshes and point clouds, such as RapidForm, MeshLab [12], and CloudCompare [13], which use rigid registration algorithms like ICP. Other solutions, such as ANTs [14], Elastix, and 3DSlicer [15], support deformable registration; however, storing and managing these transformations is significantly more complex.
@@ 3.1 Concept of the transformation tree model
$\mathbf{M^{\prime}}=\mathcal{P}\cdot\mathbf{M}=\prod_{j=J}^{1}{\mathbf{P}_{j}}\cdot\mathbf{M}=\left(\mathbf{P}_{1}\cdot\mathbf{P}_{2}\cdot\ldots\cdot\mathbf{P}_{J}\right)\cdot\mathbf{M}$ (4)
where $\mathbf{P}_{J}\cdot...\mathbf{P}_{2}\cdot\mathbf{P}_{1}$ are the transformation matrices representing subsequent transformations of the model from device to final coordinate systems, $\mathcal{P}$ — cumulative transformation matrix, $J$ — the number of subsequent transformations.
@@ 4.2 Embeding intraoral scans in 3D face scans — list of transformations
Figure 7: Transformation tree
@@ 4.4 Groups of objects — transformation and nesting
All elements of the group are subject to the same transformations.
Registration of partial 3D face scans from different viewpoints can also be performed using pairwise registration between neighbouring scans.
Each newly registered scan is aligned with the previous one, forming a group. This group is then transformed using the preceding registration, progressively expanding until all scans are iteratively registered back to the first scan. It is explained in Figure 10 and can be represented with the following equation:

### v3

@@ 2.2 Existing Approaches to Storing Image Transformations
Current methods for storing registration results vary depending on the system and standard used. Tools such as MeVisLab [9], ITK [10], and Elastix [11] provide capabilities for image registration and may return a final transformation matrix or a transformed image. However, they often lack systematic recording of the full registration process. As a result, important intermediate steps are lost, making it difficult to reconstruct the history of transformations or verify results retrospectively. Similar challenges occur in geometry-based registration tools, such as RapidForm, MeshLab [12] and CloudCompare [13], which are commonly used for registering surface meshes and point clouds using algorithms like ICP. While the data types differ, these tools share the limitation of not storing the full sequence of transformations. Additionally, deformable registration tools such as ANTs [14], Elastix [11], and 3D Slicer [15] introduce further complexity in transformation management, yet still provide limited support for comprehensive documentation.
@@ 3.1 Concept of the transformation tree model
$\mathbf{M^{\prime}}=\mathcal{P}\cdot\mathbf{M}=\prod_{j=J}^{1}{\mathbf{P}_{j}}\cdot\mathbf{M}=\left(\mathbf{P}_{J}\cdot\ldots\cdot\mathbf{P}_{2}\cdot\mathbf{P}_{1}\right)\cdot\mathbf{M}$ (4)
where $\mathbf{P}_{J}\cdot...\cdot\mathbf{P}_{2}\cdot\mathbf{P}_{1}$ are the transformation matrices representing subsequent transformations of the model from device to final coordinate systems, $\mathcal{P}$ — cumulative transformation matrix, $J$ — the number of subsequent transformations.
@@ 3.4 Summary
Hereby, indirect registration is one of the key advantages of the transformation tree. Such registration refers to the ability to determine the spatial relationship between two images that do not share a direct transformation path. In the transformation tree model, each image is linked to a common root via a unique transformation chain. Therefore, a transformation from image A to image B can be computed by composing the inverse of the path from A to the root with the path from the root to B. This allows images to be registered indirectly, without having to perform direct pairwise registration, as long as both are connected to the tree. This property is especially valuable in clinical applications involving complex, multi-stage acquisition workflows, where certain datasets cannot be registered directly due to missing reference points.
@@ 4.2 Embeding intraoral scans in 3D face scans — list of transformations
Figure 7: Example of a transformation tree representing the registration of an intraoral scan into a 3D face scan using intermediate scans with a positioner. Each edge corresponds to a transformation matrix ($P_{1}$ to $P_{4}$), and the hierarchical structure reflects the sequence of operations. The figure illustrates how multiple partial registrations are combined to form a complete transformation path within a unified coordinate system.
@@ 4.4 Groups of objects — transformation and nesting
In a transformation tree, all elements of a group share the same transformation path. This principle applies, for example, to the registration of partial 3D face scans from different viewpoints, which can be performed pairwise, forming progressively expanding groups. Each newly registered scan is aligned with the previous one, forming a group. This group is then transformed using the preceding registration, progressively expanding until all scans are iteratively registered back to the first scan. It is explained in Figure 10 and can be represented with the following equation: